OpenStudy (anonymous):

The LEngth of a Rectangle is twice its width. Its perimeter is 15 inches, what is the area of the Rectangle?

OpenStudy (anonymous):

2x (width) + 2y(length) = 15..... length is twice its width so 2x + 2(2x) = 15 2x+4x = 15 6x = 15 x = 15/6

OpenStudy (anonymous):

We have here that L=2W L being length and W being Width. Perimeter is 2L+2W so 15=2L+2W but we know that L= 2W so 15=2(2W)+2W so then 15=4W+2W or 15=6W so then W=15/6. and since L=2W L=2(15/6) = 30/6 = 5 and the area is = L* W which we now know, so all you do is multiply

OpenStudy (anonymous):

I hadn't seen that you had posted it up already -_- lol

OpenStudy (anonymous):

juan_9292 is more complete though, I didn't notice the area was required as well.

OpenStudy (anonymous):

so whats the Area then? O.o confused

OpenStudy (anonymous):

Well since by solving for W and later on for L we got that W=15/6 L=5 so for area all you have to do is Multiply L X W

OpenStudy (anonymous):

it must b wrong then my answer choices are { f:7.5, G: 12.5 , H:15, i:25.}

OpenStudy (anonymous):

no, its right;

OpenStudy (anonymous):

 which that is = 12.5